Given: x+y+10=0 and underneath it is x^2+y-2=0
What are the x-values for the solutions to the given system of equations?
Thank you for the steps to get to the solution.

Answer by stanbon(75887)

Given: x+y+10=0 and underneath it is x^2+y-2=0
What are the x-values for the solutions to the given system of equations?
-----------
x + y = -10
x^2 + y = 2
----------
Subtract and solve for "x":
x^2 - x = 12
----
x^2 - x - 12 = 0
----
Factor::
(x-4)(x+3) = 0
----
x = 4 or x = -3
------
Cheers,
Stan H.